# Optimal Frames for Phase Retrieval from Edge Vectors of Optimal Polygons

### v2

@@ Abstract.
This paper aims to characterize the optimal frame for phase retrieval, defined as the frame whose condition number for phase retrieval attains its minimal value. In the context of the two-dimensional real case, we reveal the connection between optimal frames for phase retrieval and the perimeter-maximizing isodiametric problem, originally proposed by Reinhardt in 1922. Our work establishes that every optimal solution to the perimeter-maximizing isodiametric problem inherently leads to an optimal frame in ${\mathbb{R}}^{2}$. By recasting the optimal polygons problem as one concerning the discrepancy of roots of unity, we characterize all optimal polygons. Building upon this connection, we then characterize all optimal frames with $m$ vectors in ${\mathbb{R}}^{2}$ for phase retrieval when $m\geq 3$ has an odd factor. As a key corollary, we show that the harmonic frame $E_{m}$ is not optimal for any even integer $m\geq 4$. This finding disproves a conjecture proposed by Xia, Xu, and Xu (Math. Comp., 94(356): 2931-2960). Previous work has established that the harmonic frame $E_{m}\subset{\mathbb{R}}^{2}$ is indeed optimal when $m$ is an odd integer.
Exploring the connection between phase retrieval and discrete geometry, this paper aims to illuminate advancements in phase retrieval and offer new perspectives on the perimeter-maximizing isodiametric problem.
@@ 1.1. Stability of phase retrieval
$\Phi_{\boldsymbol{A}}(\boldsymbol{x})=|\boldsymbol{A}\boldsymbol{x}|:=\left(\left|\left\langle\boldsymbol{a}_{1},\boldsymbol{x}\right\rangle\right|,\left|\left\langle\boldsymbol{a}_{2},\boldsymbol{x}\right\rangle\right|,\ldots,\left|\left\langle\boldsymbol{a}_{m},\boldsymbol{x}\right\rangle\right|\right)^{T}\in\mathbb{R}_{+}^{m}$ (1.1)
@@
A smaller value of $\beta_{\boldsymbol{A}}^{\mathbb{H}}$ entails that $\Phi_{\boldsymbol{A}}$ behaves more like a near-isometry, indicating greater stability against the measurement noise. If $\boldsymbol{A}$ lacks the phase retrieval property, then $L_{\boldsymbol{A}}^{\mathbb{H}}=0$ and $\beta_{\boldsymbol{A}}^{\mathbb{H}}=+\infty$. We say that a matrix $\boldsymbol{A}\in\mathbb{H}^{m\times d}$ has the minimal condition number if $\beta_{\boldsymbol{A}}^{\mathbb{H}}=\min_{\boldsymbol{M}\in\mathbb{H}^{m\times d}}\beta_{\boldsymbol{M}}^{\mathbb{H}}$. To simplify the notation, we often omit the superscript in $\beta_{\boldsymbol{A}}^{\mathbb{H}}$, determining whether $\beta_{\boldsymbol{A}}^{\mathbb{H}}$ is defined over the real or complex field based on whether the matrix $\boldsymbol{A}$ is real or complex. In the same way, we drop the superscripts in $L_{\boldsymbol{A}}^{\mathbb{H}},U_{\boldsymbol{A}}^{\mathbb{H}}$, and we omit the subscript in $\operatorname{dist}_{\mathbb{H}}(\boldsymbol{x},\boldsymbol{y})$ when the underlying field is clear from context.
@@
$\beta_{\boldsymbol{A}}^{\mathbb{R}}\geq\frac{1}{\sqrt{1-\frac{1}{m\cdot\sin\frac{\pi}{2m}}}}.$ (1.3)
According to Theorem 3.3 in [32], we have
$\beta_{\boldsymbol{E}_{m}}^{\mathbb{R}}=\begin{cases}\frac{1}{\sqrt{1-\frac{2}{m\cdot\sin\frac{\pi}{m}}}}&\text{if $m\geq 3$ is even,}\\ \frac{1}{\sqrt{1-\frac{1}{m\cdot\sin\frac{\pi}{2m}}}}&\text{if $m\geq 3$ is odd,}\end{cases}$ (1.4)
where $\boldsymbol{E}_{m}$ is the harmonic frame in ${\mathbb{R}}^{2}$ which is defined by
@@
Combining (1.3) and (1.4), we have $\beta_{\boldsymbol{E}_{m}}^{\mathbb{R}}=\min_{\boldsymbol{A}\in\mathbb{R}^{m\times 2}}\beta_{\boldsymbol{A}}^{\mathbb{R}}$ for each odd integer $m\geq 3$. The following conjecture is proposed in [32] for each even integer $m\geq 4$.
@@
If $m\geq 4$ is even, then
$\beta_{\boldsymbol{E}_{m}}^{\mathbb{R}}=\min_{\boldsymbol{A}\in\mathbb{R}^{m\times 2}}\beta_{\boldsymbol{A}}^{\mathbb{R}}.$
In this paper, we will show that minimizing the condition number of $\boldsymbol{A}\in\mathbb{R}^{m\times 2}$ can be transformed into solving the perimeter-maximizing isodiametric problem proposed by Reinhardt in 1922 [25]. We prove that each optimal solution to the perimeter-maximizing isodiametric problem gives rise to a matrix with the minimal condition number. Based on that, we can show that Conjecture 1.1 is false. For instance, when $m=4$, we will demonstrate in Remark 1.3 that
$\beta_{\boldsymbol{E}_{4}}^{\mathbb{R}}={\sqrt{2+\sqrt{2}}}\approx 1.84>\beta_{\boldsymbol{E}_{4}^{\prime}}^{\mathbb{R}}=\min_{\boldsymbol{M}\in\mathbb{R}^{4\times 2}}\beta_{\boldsymbol{M}}^{\mathbb{R}}=\sqrt{1+\frac{\sqrt{2}}{2}+\frac{\sqrt{6}}{2}}\approx 1.71,$
where
$\boldsymbol{E}_{4}^{\prime}=\left(\begin{array}[]{cccc}-1&\cos\frac{2\pi}{3}&\sqrt{2\sin\frac{\pi}{12}}\cos\frac{3\pi}{8}&\sqrt{2\sin\frac{\pi}{12}}\cos\frac{7\pi}{24}\\ 0&\sin\frac{2\pi}{3}&\sqrt{2\sin\frac{\pi}{12}}\sin\frac{3\pi}{8}&\sqrt{2\sin\frac{\pi}{12}}\sin\frac{7\pi}{24}\end{array}\right)^{T}.$ (1.6)
1.2. The perimeter-maximizing isodiametric problem for convex polygons
The perimeter-maximizing isodiametric problem has a long and rich history. It was first studied by Reinhardt in 1922 [25], and has been furthered investigated by Vincze [30] in 1950 and Datta [14] in 1997 (see also [15, 16, 24, 9]). The convex polygon in $\mathbb{R}^{2}$ with $m$ edges is called the convex $m$-gon. The diameter of a polygon $P$, denoted by $\operatorname{diam}({P})$, is the largest distance between any pair of its vertices, i.e., $\operatorname{diam}({P}):=\max_{\boldsymbol{x},\boldsymbol{y}\,\in{P}}{\|\boldsymbol{x}-\boldsymbol{y}\|_{2}}$. The perimeter of a polygon $P$, denoted by $\operatorname{perim}({P})$, is the sum of the length of its edges. The perimeter-maximizing isodiametric problem asks:
Problem 1.1.
Among all convex $m$-gons with fixed diameter, which ones has the maximal perimeter?
Definition 1.1.
  1. (i)
    Optimal polygon: A convex $m$-gon is defined as optimal if it is a solution to Problem 1.1. We use ${\mathcal{E}}_{m}$ to denote the collection of all optimal convex $m$-gons in $\mathbb{R}^{2}$.
  2. (ii)
  3. (iii)
  4. (iv)
  5. (v)
It is known that the optimal convex $m$-gon must be strictly convex [25]. Denote r(P)r({P}) as the diameter-to-perimeter ratio of a polygon P{P}, i.e.,
$r(P):=\frac{\mathrm{diam}(P)}{\mathrm{perim}(P)}.$ (1.7)
Then solving Problem 1.1 is equivalent to solving
$\min_{P\in\mathcal{P}_{m}}r(P).$ (1.8)
We say that the solution to (1.8) is unique if it is unique up to scaling, translations, rotations, flips, or any combination of these transformations.
Theorem 1.1 summarizes the existing results for the perimeter-maximizing isodiametric problem.
@@ 1.2. The perimeter-maximizing isodiametric problem for convex polygons
    ([25, 14, 16, 15, 30, 24]) Let $m\geq 3$ be a positive integer that has an odd factor. Then
    $\min_{{P}\in\mathcal{P}_{m}}r({P})=\frac{1}{2m\cdot\sin\frac{\pi}{2m}},$
    where the equality is achieved by finitely many equilateral strictly convex $m$-gons .
  2. (ii)
    ([25, 14, 16, 15]) Let $m=2^{s}$, where $s$ is an integer and $s\geq 2$. Then
    $\min_{{P}\in\mathcal{P}_{m}}r({P})>\frac{1}{2m\cdot\sin\frac{\pi}{2m}}.$
  3. (iii)
    ([25, 14, 16, 15, 30]) The regular $m$-gon $P_{m}$ is optimal if and only if $m$ is odd. Furthermore, ${P}_{m}$ is uniquely optimal if and only if $m$ is a prime .
1.3. Our contribution.
In this paper, we focus on estimating the minimal condition number for $m\times 2$ real matrices and characterizing all $m\times 2$ real matrices that has the minimal condition number. For convenience, we introduce the following definitions.
Definition 1.2.
A finite set in $\mathbb{R}^{d}$ is called a frame if the vectors in this set can span $\mathbb{R}^{d}$. We say that a frame $\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}\subset\mathbb{R}^{d}$ is optimal if its associated matrix $\boldsymbol{A}=[\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{m}]^{T}\in\mathbb{R}^{m\times d}$ has the minimal condition number, i.e., $\beta_{\boldsymbol{A}}^{\mathbb{R}}=\min_{\boldsymbol{M}\in\mathbb{R}^{m\times d}}\beta_{\boldsymbol{M}}^{\mathbb{R}}$. A frame $\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}\subset\mathbb{R}^{d}$ is called a tight frame if $\sum_{i=1}^{m}\boldsymbol{a}_{i}\boldsymbol{a}_{i}^{T}=C\cdot\boldsymbol{I}_{d}$ for some constant $C>0$.
Definition 1.3.
Let $f:\mathcal{T}\to\mathbb{R}^{2}$ be a map defined by
$f(\boldsymbol{a})\,\,=\,\,t^{2}\cdot(\cos 2\phi,\sin 2\phi)^{T}$ (1.10)
for any $\boldsymbol{a}=t\cdot(\cos\phi,\sin\phi)^{T}\in\mathcal{T}$.
Remark 1.1.
We will now show that the map $f:\mathcal{T}\to\mathbb{R}^{2}$ is a bijection. The bijectivity can be established by considering two cases. Firstly, for the zero vector, it holds that $\boldsymbol{a}=\boldsymbol{0}$ if and only if $f(\boldsymbol{a})=\boldsymbol{0}$. Secondly, for any nonzero vector $\boldsymbol{e}_{0}\in\mathbb{R}^{2}$, it possesses a unique representation in the form $\boldsymbol{e}_{0}=t_{0}\cdot(\cos{2\phi_{0}},\sin{2\phi_{0}})^{T}$, where $t_{0}>0$ and $\phi_{0}\in(0,\pi]$. Given this, the unique pre-image $\boldsymbol{a}_{0}\in\mathcal{T}$ such that $f(\boldsymbol{a}_{0})=\boldsymbol{e}_{0}$ is given by $\boldsymbol{a}_{0}=\sqrt{t_{0}}\cdot(\cos{\phi_{0}},\sin{\phi_{0}})^{T}$. Therefore, $f$ is both injective and surjective, implying it is a bijection. This ensures that the inverse map $f^{-1}:\mathbb{R}^{2}\to\mathcal{T}$ is well-defined, and is explicitly given by $f^{-1}(\boldsymbol{e}_{0})=\boldsymbol{a}_{0}$.
1.3.1. Estimating the minimal condition number
We first estimate the minimal condition number of real $m\times 2$ matrices. Note that the condition number of a matrix $\boldsymbol{A}=[\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{m}]^{T}\in\mathbb{R}^{m\times 2}$ does not change if the row vector $\boldsymbol{a}_{i}$ is replaced by $\boldsymbol{a}_{i}$ for any $i\in[m]:=\{1,\dotsc,m\}$. Hence, to estimate the minimal condition number, it is enough to consider the case when each $\boldsymbol{a}_{i}\in\mathcal{T}$.
We begin by establishing a bijective correspondence between optimal tight frames with $m$ vectors in $\mathbb{R}^{2}$ and optimal convex $m$-gons.
Theorem 1.2.
Assume that $\mathcal{A}=\{\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{m}\}\subset\mathcal{T}\subset{\mathbb{R}}^{2}$, where $\mathcal{T}$ is defined in (1.9). The set $\mathcal{A}$ is an optimal tight frame if and only if $\{f(\boldsymbol{a}_{1}),\dotsc,f(\boldsymbol{a}_{m})\}$ is the edge vectors set of an optimal $m$-gon $P\in\mathcal{P}_{m}$, where $f(\cdot)$ is defined in (1.10). In either case, we have
$\beta_{\boldsymbol{A}}=\frac{1}{\sqrt{1-2\cdot r({P})}},$
Combining Theorem 1.2 with Theorem 1.1, we can derive Corollary 1.1, which presents a more precise estimate on the minimal condition number for phase retrieval.
Corollary 1.1.
  1. (i)
    For any integers $m\geq 3$ which has an odd factor, we have
@@ 1.3.1. Estimating the minimal condition number
The authors in [32] proved that the lower bound (1.3) is tight if $m\geq 3$ is an odd integer. Corollary 1.1 (i) generalizes this result by showing that the lower bound (1.3) is tight as long as $m\geq 3$ has an odd factor. Moreover, Corollary 1.1 (iii) shows that Conjecture 1.1 is false, i.e., the condition number of harmonic frame $\boldsymbol{E}_{m}$ is not minimal for each even integer $m\geq 4$.
Remark 1.3.
Take $m=4$ as an example. Let $P\subset\mathbb{R}^{2}$ be the convex $4$-gon with the edge set
@@
$\left\{\begin{pmatrix}1\\ 0\end{pmatrix},\begin{pmatrix}\cos\frac{4\pi}{3}\\ \sin\frac{4\pi}{3}\end{pmatrix},2\sin\frac{\pi}{12}\begin{pmatrix}\cos\frac{3\pi}{4}\\ \sin\frac{3\pi}{4}\end{pmatrix},2\sin\frac{\pi}{12}\begin{pmatrix}\cos\frac{7\pi}{12}\\ \sin\frac{7\pi}{12}\end{pmatrix}\right\}$.
It is known that $P$ is the unique optimal $4$-gon, and we have $r(P)=\frac{1}{2+\sqrt{6}-\sqrt{2}}$ [29, 28]. Let $\boldsymbol{A}\in\mathbb{R}^{4\times 2}$ be defined as in (1.6). By Theorem 1.2 we see that the matrix $\boldsymbol{A}$ has the minimal condition number $\beta_{\boldsymbol{A}}^{\mathbb{R}}=\sqrt{1+\frac{\sqrt{2}}{2}+\frac{\sqrt{6}}{2}}\approx 1.71$. According to (1.4), we have $\beta_{\boldsymbol{E}_{4}}^{\mathbb{R}}={\sqrt{2+\sqrt{2}}}\approx 1.84$, which is strictly greater than $\beta_{\boldsymbol{A}}^{\mathbb{R}}$. This is aligned with Corollary 1.1 (iii). The optimal convex $4$-gon $P$ and the rows of $\boldsymbol{A}$ are both illustrated in Figure 1.
1.3.2. Characterization of all optimal polygons
Recall that ${\mathcal{E}}_{m}$ denotes the set of all optimal convex $m$-gons in $\mathbb{R}^{2}$. We then proceed to characterize these optimal polygons $P\in{\mathcal{E}}_{m}$. This characterization begins with the introduction of the following problem concerning the discrepancy of roots of unity. Subsequently, we will demonstrate the connection between this discrepancy problem and the optimal polygons.
Problem 1.2.
Theorem 1.3.
Let $m\geq 3$ be an integer with an odd factor. Set $\boldsymbol{\mu}_{j}:=(\cos\frac{j\pi}{m},\sin\frac{j\pi}{m})^{T}$ for each $j\in[m]$. Then we have
  1. (i)
  2. (ii)
Remark 1.4.
For a vector $\boldsymbol{\varepsilon}=(\varepsilon_{1},\dotsc,\varepsilon_{m})^{T}\in\{\pm 1\}^{m}$, we define the shifting operator $\mathcal{S}$ and the flip operator $\mathcal{R}$ as
  1. (i)
    if $P_{2}$ is obtained from $P_{1}$ through a proper translation and scaling, then $\boldsymbol{\varepsilon}_{2}=\boldsymbol{\varepsilon}_{1}$;
  2. (ii)
  3. (iii)
    if $P_{2}$ is obtained from $P_{1}$ by a flip with respect to the $y$-axis, then $\boldsymbol{\varepsilon}_{2}=\mathcal{S}(\mathcal{R}(\boldsymbol{\varepsilon}_{1}))$.
Since we treat polygons as equivalent under scaling, translations, rotations, flips, or any combination of these transformations, we define two vectors $\boldsymbol{\varepsilon}_{1},\boldsymbol{\varepsilon}_{2}\in\{\pm 1\}^{m}$ to be equivalent if one can be obtained from the other by applying a finite sequence of the shifting operator $\mathcal{S}$ and the flip operator $\mathcal{R}$. A simple calculation shows that ${\mathcal{R}}^{2}={\mathcal{S}}^{2m}={\rm Id}$, where $\rm Id$} denotes the identity operator. Furthermore, we have ${\mathcal{S}}^{k}{\mathcal{R}}={\mathcal{R}}{\mathcal{S}}^{2m-k}$ for any integer $k\in[1,2m]$. Hence, $\boldsymbol{\varepsilon}_{1}$ is equivalent to $\boldsymbol{\varepsilon}_{2}$ if and only if ${\mathcal{R}}^{k_{1}}{\mathcal{S}}^{k_{2}}(\boldsymbol{\varepsilon}_{1})=\boldsymbol{\varepsilon}_{2}$ for some $k_{1}\in\{0,1\}$ and $k_{2}\in\{0,1,\ldots,2m-1\}$. Under this convention, the number of distinct solutions to Problem 1.1 equals that of Problem 1.2.
1.3.3. Characterization of all optimal frames
@@
We next characterize all optimal frames with $m$ vectors when $m\geq 3$ has an odd factor. For convenience, for any vector ${\boldsymbol{\nu}}=t\cdot(\cos\phi,\sin\phi)^{T}\in\mathcal{T}$ with $t\geq 0$ and $\phi\in(0,\frac{\pi}{2}]$, we define ${\boldsymbol{\nu}}^{\perp}:=t\cdot(\cos(\phi+\frac{\pi}{2}),\sin(\phi+\frac{\pi}{2}))^{T}\in\mathcal{T}\subset{\mathbb{R}}^{2}$ .
@@
    If $\boldsymbol{\varepsilon}=(\varepsilon_{1},\dotsc,\varepsilon_{m})^{T}\in\{\pm 1\}^{m}$ satisfies (1.11), then the frame ${\mathcal{A}}=\{{\boldsymbol{\nu}}_{j}:j\in I\}\cup\{{\boldsymbol{\nu}}_{j}^{\perp}:j\in[m]\setminus I\}$ is an optimal frame, i.e., $\mathcal{A}\in\mathcal{T}_{m}$, where $I=\{j\in[m]:\varepsilon_{j}=1\}$.
@@
    For any optimal frame ${\mathcal{A}}\in{\mathcal{S}}_{m}$, there exists a vector $\boldsymbol{\varepsilon}=(\varepsilon_{1},\dotsc,\varepsilon_{m})^{T}\in\{\pm 1\}^{m}$ satisfying (1.11), such that ${\mathcal{A}}=\{{\boldsymbol{\nu}}_{j}:j\in I\}\cup\{{\boldsymbol{\nu}}_{j}^{\perp}:j\in[m]\setminus I\}$ where $I=\{j\in[m]:\varepsilon_{j}=1\}$, after a proper scaling and rotation.

### v3

@@ Abstract.
This paper aims to characterize the optimal frame for phase retrieval, defined as the frame whose condition number for phase retrieval attains its minimal value. In the context of the two-dimensional real case, we reveal the connection between optimal frames for phase retrieval and the perimeter-maximizing isodiametric problem, originally proposed by Reinhardt in 1922. Our work establishes that every optimal solution to the perimeter-maximizing isodiametric problem inherently leads to an optimal frame in ${\mathbb{R}}^{2}$. By recasting the optimal polygons problem as one concerning the discrepancy of roots of unity, we characterize all optimal polygons. Building upon this connection, we then characterize all optimal frames with $m$ vectors in ${\mathbb{R}}^{2}$ for phase retrieval when $m\geq 3$ has an odd factor. As a key corollary, we show that the harmonic frame $E_{m}\subset{\mathbb{R}}^{2}$ is not optimal for any even integer $m\geq 4$. This finding disproves a conjecture proposed by Xia, Xu, and Xu [Math. Comp., 94 (2025), pp. 2931–2960]. Previous work has established that $E_{m}$ is indeed optimal when $m$ is an odd integer.
@@
The second author was supported by NSFC grant (12501121, 12571105) and the Natural Science Foundation of Shanghai grant (25ZR1402131).
@@ 1.1. Problem setup
$\Phi_{\boldsymbol{A}}(\boldsymbol{x})=|\boldsymbol{A}\boldsymbol{x}|:=\left(\left|\left\langle\boldsymbol{a}_{1},\boldsymbol{x}\right\rangle\right|,\left|\left\langle\boldsymbol{a}_{2},\boldsymbol{x}\right\rangle\right|,\ldots,\left|\left\langle\boldsymbol{a}_{m},\boldsymbol{x}\right\rangle\right|\right)^{T}\in\mathbb{R}_{+}^{m}.$ (1.1)
@@
A smaller value of $\beta_{\boldsymbol{A}}^{\mathbb{H}}$ entails that $\Phi_{\boldsymbol{A}}$ behaves more like a near-isometry, indicating greater stability against measurement noise. If $\boldsymbol{A}$ lacks the phase retrieval property, then $L_{\boldsymbol{A}}^{\mathbb{H}}=0$ and $\beta_{\boldsymbol{A}}^{\mathbb{H}}=+\infty$. We say that a matrix $\boldsymbol{A}\in\mathbb{H}^{m\times d}$ has the minimal condition number if $\beta_{\boldsymbol{A}}^{\mathbb{H}}=\min_{\boldsymbol{M}\in\mathbb{H}^{m\times d}}\beta_{\boldsymbol{M}}^{\mathbb{H}}$. To simplify the notation, we often omit the superscript in $\beta_{\boldsymbol{A}}^{\mathbb{H}}$, determining whether $\beta_{\boldsymbol{A}}^{\mathbb{H}}$ is defined over the real or complex field based on whether the matrix $\boldsymbol{A}$ is real or complex. In the same way, we drop the superscripts in $L_{\boldsymbol{A}}^{\mathbb{H}},U_{\boldsymbol{A}}^{\mathbb{H}}$, and we omit the subscript in $\operatorname{dist}_{\mathbb{H}}(\boldsymbol{x},\boldsymbol{y})$ when the underlying field is clear from context.
@@
$\beta_{\boldsymbol{A}}\geq\frac{1}{\sqrt{1-\frac{1}{m\cdot\sin\frac{\pi}{2m}}}}.$ (1.3)
In particular, they provide that
$\beta_{\boldsymbol{E}_{m}}=\begin{cases}\frac{1}{\sqrt{1-\frac{2}{m\cdot\sin\frac{\pi}{m}}}}&\text{if $m\geq 3$ is even,}\\ \frac{1}{\sqrt{1-\frac{1}{m\cdot\sin\frac{\pi}{2m}}}}&\text{if $m\geq 3$ is odd,}\end{cases}$ (1.4)
where the rows of the matrix $\boldsymbol{E}_{m}$ form the harmonic frame in ${\mathbb{R}}^{2}$, i.e.,
@@
Combining (1.3) and (1.4), they further obtain that $\beta_{\boldsymbol{E}_{m}}=\min_{\boldsymbol{A}\in\mathbb{R}^{m\times 2}}\beta_{\boldsymbol{A}}$ for each odd integer $m\geq 3$. Consequently, the following conjecture is naturally proposed in [33] for each even integer $m\geq 4$.
@@
[33] If $m\geq 4$ is even, then
$\beta_{\boldsymbol{E}_{m}}=\min_{\boldsymbol{A}\in\mathbb{R}^{m\times 2}}\beta_{\boldsymbol{A}}.$
1.2. Our contribution
In this paper, we will disprove Conjecture 1.1 by deriving a sharper estimate for the minimal condition number of matrices $\boldsymbol{A}\in\mathbb{R}^{m\times 2}$. Our key observation is that minimizing the condition number of $\boldsymbol{A}\in\mathbb{R}^{m\times 2}$ is equivalent to solving a classical problem in discrete geometry called the perimeter-maximizing isodiametric problem, which was introduced by Reinhardt in 1922 [26]. This equivalence leads directly to an improved bound on the minimal condition number, which we state in the following theorem.
@@ 1.2. Our contribution
    For any integer $m\geq 3$ that has an odd factor, we have
@@
The authors in [33] proved that the lower bound (1.3) is tight if $m\geq 3$ is an odd integer. Theorem 1.1 (i) generalizes this result by showing that the lower bound (1.3) is tight as long as $m\geq 3$ has an odd factor. Moreover, Theorem 1.1 (iii) shows that Conjecture 1.1 is false. For instance, when $m=4$, we will demonstrate in Remark 3.2 that
$\beta_{\boldsymbol{A}}=\min_{\boldsymbol{M}\in\mathbb{R}^{4\times 2}}\beta_{\boldsymbol{M}}=\sqrt{1+\frac{\sqrt{2}}{2}+\frac{\sqrt{6}}{2}}\approx 1.71,$
where
@@
$\boldsymbol{A}=\left(\begin{array}[]{cccc}-1&\cos\frac{2\pi}{3}&\sqrt{2\sin\frac{\pi}{12}}\cos\frac{3\pi}{8}&\sqrt{2\sin\frac{\pi}{12}}\cos\frac{7\pi}{24}\\ 0&\sin\frac{2\pi}{3}&\sqrt{2\sin\frac{\pi}{12}}\sin\frac{3\pi}{8}&\sqrt{2\sin\frac{\pi}{12}}\sin\frac{7\pi}{24}\end{array}\right)^{T}.$ (1.6)
In contrast, by (1.4) we have $\beta_{\boldsymbol{E}_{4}}={\sqrt{2+\sqrt{2}}}\approx 1.84$, showing that $\beta_{\boldsymbol{E}_{4}}$ is not minimal.
We next characterize all matrices in $\mathbb{R}^{m\times 2}$ with the minimal condition number, where $m\geq 3$ has an odd factor. For convenience, we introduce the following definitions.
Definition 1.1.
A finite set in $\mathbb{R}^{d}$ is called a frame if it spans $\mathbb{R}^{d}$. We say that a frame $\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}\subset\mathbb{R}^{d}$ is optimal if its associated matrix $\boldsymbol{A}=[\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{m}]^{T}\in\mathbb{R}^{m\times d}$ has the minimal condition number, i.e., $\beta_{\boldsymbol{A}}=\min_{\boldsymbol{M}\in\mathbb{R}^{m\times d}}\beta_{\boldsymbol{M}}$. A frame $\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}\subset\mathbb{R}^{d}$ is called a tight frame if $\sum_{i=1}^{m}\boldsymbol{a}_{i}\boldsymbol{a}_{i}^{T}=C\cdot\boldsymbol{I}_{d}$ for some constant $C>0$.
Definition 1.2.
@@
For any vector ${\boldsymbol{\nu}}=t\cdot(\cos\phi,\sin\phi)^{T}\in\mathcal{T}$ with $t\geq 0$ and $\phi\in(0,\frac{\pi}{2}]$, we define ${\boldsymbol{\nu}}^{\perp}:=t\cdot(\cos(\phi+\frac{\pi}{2}),\sin(\phi+\frac{\pi}{2}))^{T}\in\mathcal{T}$.
Note that the condition number of a matrix $\boldsymbol{A}=[\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{m}]^{T}\in\mathbb{R}^{m\times 2}$ does not change if the row vector $\boldsymbol{a}_{i}$ is replaced by $\boldsymbol{a}_{i}$ for any $i\in[m]:=\{1,\dotsc,m\}$. Hence, to characterize all optimal frames in $\mathbb{R}^{2}$, it is enough to consider the case when each $\boldsymbol{a}_{i}\in\mathcal{T}$.
@@
    If $\boldsymbol{\varepsilon}=(\varepsilon_{1},\dotsc,\varepsilon_{m})^{T}\in\{\pm 1\}^{m}$ satisfies
    $\sum_{j=1}^{m}\varepsilon_{j}\zeta_{m}^{j}=0,$ (1.8)
    where $\zeta_{m}:=e^{{\rm i}\pi/m}$, then the frame ${\mathcal{A}}=\{{\boldsymbol{\nu}}_{j}:j\in I\}\cup\{{\boldsymbol{\nu}}_{j}^{\perp}:j\in[m]\setminus I\}$ is an optimal frame, i.e., $\mathcal{A}\in\mathcal{T}_{m}$, where $I=\{j\in[m]:\varepsilon_{j}=1\}$.
@@
    For any optimal frame ${\mathcal{A}}\in{\mathcal{T}}_{m}$, there exists a vector $\boldsymbol{\varepsilon}=(\varepsilon_{1},\dotsc,\varepsilon_{m})^{T}\in\{\pm 1\}^{m}$ satisfying (1.8), such that ${\mathcal{A}}=\{{\boldsymbol{\nu}}_{j}:j\in I\}\cup\{{\boldsymbol{\nu}}_{j}^{\perp}:j\in[m]\setminus I\}$ where $I=\{j\in[m]:\varepsilon_{j}=1\}$, after a proper scaling and rotation.